# Context parroting: A simple but tough-to-beat baseline for foundation models in scientific machine learning

### v1

@@ C.1 Discrete-token parroting
For fully-discrete tokens, an $L^{th}$ order Markov chain fit to the context has the form
$p(\mathbf{y}|\mathbf{q})=\dfrac{\#\{j:(x_{j-(L-1):j}=\mathbf{q})\land(x_{j+1:j+H}=\mathbf{y})}{\sum_{y^{\prime}}\#\{i:(x_{j-(L-1):j}=\mathbf{q})\land(x_{j+1:j+H}=\mathbf{y^{\prime}})}$ (3)
@@
where the overall context has length $C$, and the Markov chain conditions the $H<C$ future tokens on the $L<C$ preceding tokens. The index $j$ runs over all contiguous sequences of length $L+H$ in the context, $j\in\{L-1,L,...,C-H-2,C-H-1\}$. The vector $\mathbf{q}\in\mathbb{R}^{L}$ represents the query, and the vector $\mathbf{y}\in\mathbb{R}^{H}$ represents the prediction in response to this query. Eq. 3 simply counts the number of token sequences of length $H+L$ that start with a given sequence of $L$ query tokens. A maximum-likelihood estimator derived from this model always samples the highest-likelihood sequence $\mathbf{y}$,
@@
$p(\mathbf{y}|\mathbf{q})=\dfrac{\#\{j:(x_{j-(L-1):j}=\mathbf{q})\land(x_{j+1:j+H}=\mathbf{y})+\alpha}{\sum_{\mathbf{y}^{\prime}}\left(\#\{i:(x_{j-(L-1):j}=\mathbf{q})\land(x_{j+1:j+H}=\mathbf{y^{\prime}})+\alpha\right)}$ (4)
@@
C.2 Continuous-token parroting
A more general time series model treats tokens as continuous-valued. Some time series foundation models like Chronos use quantile binning to discretize time series values, allowing the direct use of discrete-token architectures [35]. However, many time series models assume effective continuity in token values, and we favor a continuous formulation in order to highlight connections to dynamical systems theory.
To model continuous-valued tokens directly, we replace the discrete count in §C.1 with a kernel-weighted estimate over all past subsequences. Let $\{\mathbf{x}_{t}\}_{t=1}^{C}$ be a univariate or multivariate time series. For context length $L$ and prediction horizon $H$, the Nadaraya–Watson estimate of the conditional density is
$\hat{p}(\mathbf{y}\mid\mathbf{q})=\frac{\sum_{j=L}^{C-H}K_{h}\left(\mathbf{q},\,\mathbf{x}_{j-(L-1):j}\right)\;K_{h}\left(\mathbf{y},\,\mathbf{x}_{j+1:j+H}\right)}{\sum_{j=L}^{C-H}K_{h}\left(\mathbf{q},\,\mathbf{x}_{j-(L-1):j}\right)},$ (5)
where $K_{h}(\mathbf{u},\mathbf{v})=h^{-d}K\bigl{(}(\mathbf{u}-\mathbf{v})/h\bigr{)}$ is a symmetric kernel with bandwidth $h$ in dimension $d=L\cdot\dim(x_{t})$. Assuming mean-squared error as a distance function in sequence space, we use a Gaussian kernel
@@ C.2 Continuous-token parroting
In practice, we drop the second kernel on $\mathbf{y}$ in Eq. 5 and write the conditional mean predictor
$\hat{\mathbf{y}}(\mathbf{q})=\sum_{j=L}^{C-H}w(\mathbf{q},\mathbf{x}_{j-(L-1):j})\;\mathbf{x}_{j+1:j+H}$ (6)
@@
$w(\mathbf{q},\mathbf{z})\equiv\frac{K_{h}\left(\mathbf{q},\,\mathbf{z}\right)}{\sum_{j=L}^{C-H}K_{h}\left(\mathbf{q},\,\mathbf{x}_{j-(L-1):j}\right)}.$
Nearest-neighbor and global average limits. The bandwidth $h$ plays the role of a smoothing parameter (analogous to $\alpha$ in Eq. 4). As $h\to 0$ the scheme approximates a single-nearest neighbor parrot, while as $h\to\infty$ it converges to a global average over all sequences.
@@
then Eq. 6 recovers a simplified form of softmax-attention, with the temperature hyperparameter $\tau$ controlling smoothness. In this view, the continuous parroting scheme is a kernel-regression analogue of discrete $k$–gram smoothing [85].
k-nearest-neighbor limit. We define a set $\text{Top}k$ corresponding to a subset of the possible values of the index $j\in\{L,L+1,...,C-H-1,C-H\}$. The $k$ elements of $\text{Top}k$ correspond to the indices $j$ that produce the $k$ largest values of $w(\mathbf{q},\mathbf{x}_{j-(L-1):j})$ across all values of $j$. We compute a simple average of these $k$ closest matches
$\hat{\mathbf{y}}(\mathbf{q})=\frac{1}{k}\sum_{j\in\text{Top}k}w(\mathbf{q},\mathbf{x}_{j-(L-1):j})\;\mathbf{x}_{j+1:j+H}$ (7)
yielding a $k$–nearest-neighbors parroting scheme. As $k$ increases, this estimator interpolates between exact parroting ($k=1$) and global average ($k\to C$).
@@
Simplex projection, a classical forecasting method in nonlinear dynamics, corresponds to the special case $H=1$ (single step prediction), $k=L+1$ in Eq. 7. The condition $k=L+1$ represents the minimal number of affinely independent neighbors needed to triangulate a point in an $L$-dimensional space [71].
@@
where the scale parameter $\bar{d}$ is determined by the distribution of distances among queries and points in the context. In practice, this parameter is often set to the mean pairwise distance among all sequences of length $L$ in the context. The optimal value of the hyperparameter $\theta$ increases as the underlying dynamics become more strongly nonlinear [72]. We note that, in the classical formulation of the S-map, a locally-linear model is fit based on all sequences of length $L+H$ seen in the context, while here we use the Nadaraya–Watson estimator in order to emphasize connections to modern kernel regression.
C.3 Invariants of motion
@@
[…]
@@
$\lim_{C\to\infty}\mathbb{E}_{p}\!\left[F(\mathbf{y})|\mathbf{q}\right]=\mathbb{E}_{\mu}\!\left[F(\mathbf{x})\right]$
where $C$ is the context length for a Nadaraya–Watson estimator $p$, $F(\mathbf{y})$ is an estimate on a sequence $\mathbf{y}$ of an invariant property of an ergodic dynamical system with measure $\mu$, and $\mathbf{q}$ is an arbitrary sequence of consecutive timepoints from the dynamical system. This proposition states that, when the context is sufficiently long, a Nadaraya–Watson estimator of an ergodic system preserves the invariant values of the underlying dynamics.
@@
[…]
We multiply both the numerator and denominator by $1/C$ and take the limit $C\to\infty$, in order to convert the summations to expectations,
[…]
where $\mathbf{x}_{\leftarrow}$ represents the first $L$ points of random lookback window of length $L+H$ sampled from the underlying dynamical system, while $\mathbf{x}_{\rightarrow}$ denotes the next $H$ timepoints. In practice, this corresponds to a time series of $L+H$ points generated by simulating the dynamics starting at a point on the attractor randomly-sampled according to the measure $\mu$.
If we take the limit $h\to 0$, then the kernel $K_{h}$ becomes a delta function, yielding
$\lim_{h\to 0}\lim_{C\to\infty}\mathbb{E}_{\mu}\!\left[F(\mathbf{y})|\mathbf{q}\right]=\mathbb{E}_{\mu}\!\left[F(\mathbf{x}_{\rightarrow})|\mathbf{x}_{\leftarrow}=\mathbf{q}\right]$
@@
$\lim_{h\to 0}\lim_{C\to\infty}\mathbb{E}_{\mu}\!\left[F(\mathbf{y})|\mathbf{q}\right]=\mathbb{E}_{\mu}\!\left[F(\mathbf{x})\right]$
C.4 Scaling laws
For a time series $\mathbf{x}_{1:T}$ with autocorrelation given by
@@ C.4 Scaling laws
the expected forecast error scales as

### v2

@@ F.1 Discrete-token parroting
For fully-discrete tokens, a $D^{th}$ order Markov chain fit to the context has the form
$p(\mathbf{y}|\mathbf{q})=\dfrac{\#\{j:(x_{j-(D-1):j}=\mathbf{q})\land(x_{j+1:j+H}=\mathbf{y})}{\sum_{y^{\prime}}\#\{i:(x_{j-(D-1):j}=\mathbf{q})\land(x_{j+1:j+H}=\mathbf{y^{\prime}})}$ (3)
@@
where the overall context has length $L$, and the Markov chain conditions the $H<L$ future tokens on the $D<L$ preceding tokens. The index $j$ runs over all contiguous sequences of length $D+H$ in the context, $j\in\{D-1,D,...,L-H-2,L-H-1\}$. The vector $\mathbf{q}\in\mathbb{R}^{D}$ represents the query, and the vector $\mathbf{y}\in\mathbb{R}^{H}$ represents the prediction in response to this query. Eq. 3 simply counts the number of token sequences of length $D+H$ that start with a given sequence of $D$ query tokens. A maximum-likelihood estimator derived from this model always samples the highest-likelihood sequence $\mathbf{y}$,
@@
$p(\mathbf{y}|\mathbf{q})=\dfrac{\#\{j:(x_{j-(D-1):j}=\mathbf{q})\land(x_{j+1:j+H}=\mathbf{y})+\alpha}{\sum_{\mathbf{y}^{\prime}}\left(\#\{i:(x_{j-(D-1):j}=\mathbf{q})\land(x_{j+1:j+H}=\mathbf{y^{\prime}})+\alpha\right)}$ (4)
@@
F.2 Continuous-token parroting
A more general time series model treats tokens as continuous-valued. Some time series foundation models like Chronos use binning to discretize time series values, allowing the direct use of discrete-token architectures [35]. However, many time series models assume effective continuity in token values, and we favor a continuous formulation in order to highlight connections to dynamical systems theory.
To model continuous‐valued tokens directly, we replace the discrete count in §F.1 with a kernel‐weighted estimate over all past subsequences. Let $\{\mathbf{x}_{t}\}$ denote a univariate or multivariate time series. For context length $L$ and prediction horizon $H$, the Nadaraya–Watson estimate of the conditional density is
$\hat{p}(\mathbf{y}\mid\mathbf{q})=\frac{\sum_{j=D}^{L-H}K_{h}\left(\mathbf{q},\,\mathbf{x}_{j-(D-1):j}\right)\;K_{h}\left(\mathbf{y},\,\mathbf{x}_{j+1:j+H}\right)}{\sum_{j=D}^{L-H}K_{h}\left(\mathbf{q},\,\mathbf{x}_{j-(D-1):j}\right)},$ (5)
where $K_{h}(\mathbf{u},\mathbf{v})=h^{-d}K\bigl((\mathbf{u}-\mathbf{v})/h\bigr)$ is a symmetric kernel with bandwidth $h$ in dimension $d=D\cdot\dim(\mathbf{x}_{t})$ for the first kernel, and $d=H\cdot\dim(\mathbf{x}_{t})$ for the second kernel. Assuming mean-squared error as a distance function in sequence space, we use a Gaussian kernel
@@ F.2 Continuous-token parroting
In practice, we drop the second kernel on $\mathbf{y}$ in Eq. 5 in order to output a prediction that exactly matches sequences from the context, rather than nearby sequences in a least-squares sense. We thus write the conditional mean predictor
$\hat{\mathbf{y}}(\mathbf{q})=\sum_{j=D}^{L-H}w(\mathbf{q},\mathbf{x}_{j-(D-1):j})\;\mathbf{x}_{j+1:j+H}$ (6)
@@
$w(\mathbf{q},\mathbf{z})\equiv\frac{K_{h}\left(\mathbf{q},\,\mathbf{z}\right)}{\sum_{j=D}^{L-H}K_{h}\left(\mathbf{q},\,\mathbf{x}_{j-(D-1):j}\right)}.$
Nearest‐neighbor and global average limits. The bandwidth $h$ plays the role of a smoothing parameter (analogous to $\alpha$ in Eq. 4). As $h\to 0$ the scheme approximates a single-nearest neighbor parrot, while as $h\to\infty$ it converges to a global average over all sequences.
@@
then Eq. 6 recovers a simplified form of softmax‐attention, with the temperature hyperparameter $\tau$ controlling smoothness. In this view, the continuous parroting scheme is a kernel‐regression analogue of discrete kk–gram smoothing [93].
k-nearest‐neighbor limit. We define a set $\text{Top}k$ corresponding to a subset of the possible values of the index $j\in\{D,D+1,...,L-H-1,L-H\}$. The $k$ elements of $\text{Top}k$ correspond to the indices $j$ that produce the $k$ largest values of $w(\mathbf{q},\mathbf{x}_{j-(D-1):j})$ across all values of $j$. We compute a simple average of these $k$ closest matches
$\hat{\mathbf{y}}(\mathbf{q})=\frac{1}{k}\sum_{j\in\text{Top}k}w(\mathbf{q},\mathbf{x}_{j-(D-1):j})\;\mathbf{x}_{j+1:j+H}$ (7)
yielding a kk–nearest-neighbors parroting scheme. As $k$ increases, this estimator interpolates between exact parroting ($k=1$) and global average ($k\to L$).
@@
Simplex projection, a classical forecasting method in nonlinear dynamics, corresponds to the special case $H=1$ (single step prediction), $k=D+1$ in Eq. 7. The condition $k=D+1$ represents the minimal number of affinely independent neighbors needed to triangulate a point in a $D$-dimensional space [74].
@@
where the scale parameter $\bar{d}$ is determined by the distribution of distances among queries and points in the context. In practice, this parameter is often set to the mean pairwise distance among all sequences of length $D$ in the context. The optimal value of the hyperparameter $\theta$ increases as the underlying dynamics become more strongly nonlinear [75]. We note that, in the classical formulation of the S-map, a locally-linear model is fit based on all sequences of length $D+H$ seen in the context, while here we use the Nadaraya–Watson estimator in order to emphasize connections to modern kernel regression.
F.3 Invariants of motion
@@
$\mathbb{E}_{\mu}\!\left[F\right]\equiv\lim_{T\to\infty}\dfrac{1}{T}\int_{0}^{T}F(\mathbf{x})=\int F(\mathbf{x})d\mu(\mathbf{x})=\text{constant}$
@@
where $L$ is the context length for a Nadaraya–Watson estimator $p$, $F(\mathbf{y})$ is an estimate on a sequence $\mathbf{y}$ of an invariant property of an ergodic dynamical system with measure $\mu$, and $\mathbf{q}$ is an arbitrary sequence of consecutive timepoints from the dynamical system. This proposition states that, when the context is sufficiently long, a Nadaraya–Watson estimator of an ergodic system preserves the invariant values of the underlying dynamics.
@@
$\mathbb{E}_{\mu}\!\left[F(\mathbf{y})|\mathbf{q}\right]=\frac{\sum_{j=D}^{L-H}K_{h}\left(\mathbf{q},\,\mathbf{x}_{j-(D-1):j}\right)\;\int F(\mathbf{y})K_{h}\left(\mathbf{y},\,\mathbf{x}_{j+1:j+H}\right)d\mu(\mathbf{y})}{\sum_{j=D}^{L-H}K_{h}\left(\mathbf{q},\,\mathbf{x}_{j-(D-1):j}\right)},$
We multiply both the numerator and denominator by $1/L$ and take the limit $L\to\infty$, in order to convert the summations to expectations,
$\lim_{L\to\infty}\mathbb{E}_{\mu}\!\left[F(\mathbf{y})|\mathbf{q}\right]=\frac{\mathbb{E}_{\mu}\!\left[K_{h}(\mathbf{q},\mathbf{x}_{\leftarrow})\int F(\mathbf{y})K_{h}\left(\mathbf{y},\mathbf{x}_{\rightarrow}\right)d\mu(\mathbf{y})\right]}{\mathbb{E}_{\mu}\!\left[K_{h}(\mathbf{q},\mathbf{x}_{\leftarrow})\right]},$
where $\mathbf{x}_{\leftarrow}$ represents the first $D$ points of random lookback window of length $D+H$ sampled from the underlying dynamical system, while $\mathbf{x}_{\rightarrow}$ denotes the next $H$ timepoints. In practice, this corresponds to a time series of $D+H$ points generated by simulating the dynamics starting at a point on the attractor randomly-sampled according to the measure $\mu$.
If we take the limit $h\to 0$ (exact matching), then the kernel $K_{h}$ becomes a delta function, yielding
$\lim_{h\to 0}\lim_{L\to\infty}\mathbb{E}_{\mu}\!\left[F(\mathbf{y})|\mathbf{q}\right]=\mathbb{E}_{\mu}\!\left[F(\mathbf{x}_{\rightarrow})|\mathbf{x}_{\leftarrow}=\mathbf{q}\right]$
@@
$\lim_{h\to 0}\lim_{L\to\infty}\mathbb{E}_{\mu}\!\left[F(\mathbf{y})|\mathbf{q}\right]=\mathbb{E}_{\mu}\!\left[F(\mathbf{x})\right]$
F.4 Scaling laws limiting prediction of stochastic systems
For a stochastic time series $\mathbf{x}_{1:T}$ with autocorrelation given by
@@ F.4 Scaling laws limiting prediction of stochastic systems
with $C$ representing a proportionality constant, the expected mean squared error of a forecast scales as